OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

I know that I need to find the slope, in fact, I've already done that..

OpenStudy (anonymous):

but how do I find the y-intercept?

OpenStudy (anonymous):

y intercept 20 slope -3 http://www.wolframalpha.com/input/?i=find+equation+%286%2C+2%29%2C+%288%2C+-4%29

OpenStudy (anonymous):

Hmm you dont need to find the intercept. First calculate the gradient (skip this if you already have it) Gradient is -3 ( skipping working) SInce the line equation is y=mx+ c where m is the gradient and c is a constant. Substitute one of the points in ( letz say P(6,2) ) and gradient= -3 into the equation. 2= (-3)6 + c c= 20 SO eqn, of line is y=-3x + 20

OpenStudy (anonymous):

so, If I were to fill in this blank _ x _ y _

OpenStudy (anonymous):

x=20/3 y=20

OpenStudy (anonymous):

what would I need to do? @tiaph that seems to make some sense, but I'm not sure.

OpenStudy (anonymous):

oh yea silly me that is in fact calculating the y-intercept, discount what i said in the first line.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hmm what do you mean what you have to do? i shown you the workings? which part you don't understand?

OpenStudy (anonymous):

Well I have one blank before x and one after that, which is the one before y _ x _ y

OpenStudy (anonymous):

go to site http://www.wolframalpha.com/input/?i=find+equation+%286%2C+2%29%2C+%288%2C+-4%29 every step is given there

OpenStudy (anonymous):

whr did you get the blanks?

OpenStudy (anonymous):

From school, it is a fill in the blank :)

OpenStudy (anonymous):

OIC, because in standard form of an equation of a line is given in the form Ax+By = C. Note that previously i said y=mx+c is an alternative version, both are right though. So in your case Ax+By = C. Compared to the answer i give, y=-3x + 20 Rearrange to get, 3x+y=20 So A =3 B = 1 C=20

OpenStudy (anonymous):

Oh, I see, thank you!

OpenStudy (anonymous):

Np.